OpenStudy (sgayton27):

Evaluate the expression without a calculator, write answer in radians: arctan(-sqrt3/3)

OpenStudy (sgayton27):

\[\tan^{-1} (-\sqrt{3}/3)\]

OpenStudy (sgayton27):

\[\tan \theta= -\sqrt{3}/3\]

OpenStudy (sgayton27):

I dont understand where tan is negative sqrt 3/3